OpenStudy (anonymous):

integral of rad(1+sin2x+(cos2x/rad(1+sin2x))^2) 0

OpenStudy (psymon):

\[\int\limits_{}^{}\sqrt{1+\sin2x+(\frac{ \cos2x }{ \sqrt{1+\sin2x} })^{2}}dx\] Hard to tell if that was what you meant or not.

OpenStudy (anonymous):

o my bad its 0<x<pi*rad2

OpenStudy (anonymous):

and yes that is what i mean

OpenStudy (psymon):

Alright, I got it, lol

OpenStudy (psymon):

So that 2nd power distributes and gives us: \[\int\limits_{}^{}\sqrt{1+\sin2x+\frac{ \cos ^{2}2x }{ 1+\sin2x }}dx\] Now that cos^2 Im going to change using the pythagorean identity: \[\int\limits_{}^{}\sqrt{1+\sin2x+\frac{ 1-\sin ^{2}2x }{ 1+\sin2x }}dx\] Now I can take that 1 - sin^2 and factor it like a difference of squares: \[\int\limits_{}^{}\sqrt{1+\sin2x+\frac{ (1-\sin2x)(1+\sin2x) }{ 1+\sin2x }}dx\] Im sure you can get it from there : )

OpenStudy (anonymous):

thx again! :D

OpenStudy (psymon):

Lol, sure xD